### v1

# LLoCO: Learning Long Contexts Offline

Abstract
Processing long contexts remains a challenge for large language models (LLMs) due to the quadratic computational and memory overhead of the self-attention mechanism and the substantial KV cache sizes during generation. We propose a novel approach to address this problem by learning contexts offline through context compression and in-domain parameter-efficient finetuning. Our method enables an LLM to create a concise representation of the original context and efficiently retrieve relevant information to answer questions accurately. We introduce LLoCO, a technique that combines context compression, retrieval, and parameter-efficient finetuning using LoRA. Our approach extends the effective context window of a 4k token LLaMA2-7B model to handle up to 128k tokens. We evaluate our approach on several long-context question-answering datasets, demonstrating that LLoCO significantly outperforms in-context learning while using 30×30$\times$30 × fewer tokens during inference. LLoCO achieves up to 7.62×7.62$\times$7.62 × speed-up and substantially reduces the cost of long document question answering, making it a promising solution for efficient long context processing. Our code is publicly available at https://github.com/jeffreysijuntan/lloco.
**footnotetext: Equal contribution
1 Introduction
Large language models (LLMs) have demonstrated remarkable performance across a wide range of tasks (Touvron et al., 2023; Jiang et al., 2023a). Many of these tasks require LLMs to understand and reason about long contexts. For instance, document question answering is one of the most common applications of LLMs, where the model is presented with a document as context and asked to respond to related questions or summarize the text. These documents may range from lengthy articles to entire books, potentially surpassing the context window limit of LLMs. Consequently, there is a growing trend in both academia and industry to enhance LLMs’ capability to process longer contexts effectively (Chen et al., 2023b; Jiang et al., 2023a; Peng et al., 2024; Chen et al., 2024). This need has driven innovation among LLM providers like OpenAI and Anthropic to develop models that can handle increasingly lengthy texts consisting of several thousands of tokens.
Despite the impressive progress made by the LLM model providers, scaling these models to adeptly manage extended contexts remains a formidable challenge, both technically and financially. Due to the self-attention mechanism, transformer-based LLMs incur a quadratic computational and memory overhead as sequence length increases. Many long-context tasks require repeatedly processing over the same prompt, which incurs extra latency overhead and substantial costs, as most commercial LLMs operate on a pricing model that is directly tied to the number of tokens processed. For example, a single inference run with a document that has 100k tokens would take 1.5 USD on Claude 3 Opus [^1] https://www.anthropic.com/api and 1 USD on GPT-4-turbo [^2] https://openai.com/pricing.
Orthogonal to the exiting efforts of expanding the context window limit, our study introduces an innovative strategy to tackle the long context challenge. We propose a method where context information is condensed offline through finetuning, enabling the model to provide accurate responses during inference with streamlined contextual representations. To illustrate our idea, consider an analogy: envision an LLM as a student preparing for an exam, where we, the researchers, are the examiners providing study materials and questions. Traditional in-context learning with full context or Retrieval-Augmented Generation (RAG) resembles an open-book exam, where the LLM has access to all materials while answering questions. In contrast, our approach is akin to a semi-closed-book exam, where the LLM cannot bring the entire book but is allowed to bring a cheat sheet. To excel in the exam, the student must 1) study efficiently to distill a concise yet informative cheat sheet, and 2) effectively retrieve relevant information from the cheat sheet to accurately answer exam questions. Namely, 1) How can we train a model to produce a compact representation of the original context that the LLM can interpret and utilize effectively? 2) How to enable the LLM to proficiently navigate and extract pertinent details from this representation during inference?
Regarding the first question, we find the formulation closely resembles the existing research direction on context compression. Prior works (Chevalier et al., 2023; Mu et al., 2023; Ge et al., 2023) have proposed compressors that aim to distill the essence of the original texts into compact representations that are aligned with LLMs. Our work primarily focuses on the second question – we’ve observed that despite the progress in context compression, LLMs often struggle to accurately read such “cheat sheets” and tend to hallucinate when applying them to answer queries. To address this issue, we employ in-domain parameter-efficient finetuning directly on the compacted context (cheat sheet) without altering its content, which significantly improves the LLM’s ability to accurately extract and utilize information from these compressed representations. In this way, we can steer the LLM to process long context more efficiently and accurately. This approach allows us to extend the effective context window of a 4k LLaMA2-7B model to handle up to 128k tokens. Moreover, we achieve state-of-the-art results that match or even surpass the performance of a LLaMA2-7B-32k model with full context on long context benchmarks, while using 30 times fewer tokens.
Figure 1: The architecture of regular LLM (left) vs LLoCO (right). In regular LLM, the long context gets appended to the prompt directly. In LLoCO, the long context goes through the context encoder first. The output summary token embeddings, which get prepended to the LLM’s prompt, are significantly shorter than the original token embeddings.
This insight has led us to introduce LLoCO, a pipeline that learns contexts offline through the combination of context compression and parameter-efficient finetuning. Our pipeline consists of three stages: preprocessing, finetuning, and serving. First, we preprocess the documents into “cheat sheets”. Then, we employ Low-Rank Adaptation (LoRA) (Hu et al., 2022) to perform parameter-efficient finetuning on these “cheat sheets” in groups. During model serving, we use a standard RAG retriever to retrieve the compressed document as well as the most relevant LoRA module, and apply them to the LLM for inference. The contributions of our work could be summarized as follows:
  • We investigate how to learn long contexts effectively and demonstrate that through a novel combination of context compression and instruction finetuning, we can extend the context window of a 4k LLaMA2-7B to handle context lengths of up to 128k tokens, while achieving performance that greatly surpasses in-context learning with 30 times fewer tokens and 8.34×8.34$\times$8.34 × latency speed-up.
  • We propose LLoCO, a novel pipeline that combines context compression, retrieval, and parameter-efficient finetuning. This pipeline could be deployed to significantly speed up and reduce the cost of long document question answering.
2 Related work
Long-context LLMs Recently, there have been efforts to increase the LLMs’ context window sizes efficiently with continual pretraining or finetuning. One line of work focuses on scaling the Rotary Position Embeddings (RoPE) (Su et al., 2021), achieving longer contexts up to $128$(Chen et al., 2023b; 2024; Peng et al., 2024). Mistral (Jiang et al., 2023a) proposes sliding window attention that only attends to part of tokens from the previous layer, reducing compute and enabling pretraining with long-context to $30$k. Nevertheless, as the auto-regressive generation in LLMs is largely memory-bound (Kwon et al., 2023), storing the KV cache of longer contexts slows down the inference and requires large GPU VRAMs.
Context compression A closely related topic is context compression, which aims to train a general compressor that can compress any input prompts. GIST (Mu et al., 2023), AutoCompressor (Chevalier et al., 2023), and ICAE (Ge et al., 2023) finetune LLMs in a “soft prompt-tuning” manner, either applying specific regularization in attention masks or utilizing dedicated “memory tokens” to compress contexts into embeddings with significant shorter lengths. LLMLingua family (Jiang et al., 2023c; b; Pan et al., 2024) proposes a question-aware framework to compress prompts in natural languages for black-box LLM APIs. Another line of work employs KV cache compression by eviction (Zhang et al., 2024b; Xiao et al., 2024), which only keeps informative keys and values for generation during inference, or quantization (Sheng et al., 2023b; Hooper et al., 2024). However, all those approaches aim to compress any inputs, which usually incurs rapid performance drops as the compression ratio exceeds a certain limit (e.g. $2$-4×4$\times$4 ×), especially for out-of-distribution texts, hindering their practical applications. In this work, we mainly focus on extreme compression of in-distribution documents of interest, going up to 30×30$\times$30 × compression rates.
Retrieval-augmented Generation Retrieval enhances the performance of LLMs in knowledge-intensive tasks such as question answering with long documents or in open-domain. RAG techniques are effective both in finetuned scenarios and when applied to off-the-shelf LLMs (Guu et al., 2020; Lewis et al., 2020; Zhang et al., 2024a), and have seen many advancements recently, emphasizing improved external knowledge integration, broader query support, and enhanced content clarity (Jiang et al., 2023d; Schick et al., 2023; Asai et al., 2023). Despite the improved performance from retrieval, challenges still remain in terms of runtime efficiency (Mallen et al., 2022) and effective filtering of irrelevant parts (Shi et al., 2023; Xu et al., 2024). Our proposed method opens up new opportunities for capturing important information while ensuring efficient generation.
Parameter-efficient Finetuning Parameter-efficient finetuning methods (Hu et al., 2022; Lester et al., 2021; Liu et al., 2024b) freeze the majority of the model weights and only update a small subset to finetune large models efficiently. LoRA (Hu et al., 2022) is one of the most widely adopted techniques, and recent advancements (Sheng et al., 2023a; Chen et al., 2023a) have focused on enhancing the system performance for deploying LoRA adaptors. In particular, Sheng et al. (2023a) has achieved the deployment of thousands of LoRA instances on a single Nvidia A100 GPU. LLoCO, which utilizes LoRA for finetuning, could benefit from those improvements to deploy LoRA adaptors efficiently in long context scenarios.
3 LLoCO
3.1 Architecture Overview
Figure 1 illustrates our proposed architecture, which consists of two components: a context encoder and an LLM decoder. In a typical instruction-finetuned LLM, the prompts can be categorized into system, user, and assistant prompts. The system prompt contains task instructions and rules the LLM should follow, as well as relevant contexts. The user prompt is a question asked by the user, and the assistant prompt is the answer generated by the LLM. The context information in the system prompt can include the user’s interaction history or documents related to the user’s question. These contexts can become very long, potentially surpassing the LLM’s context window limit.
To overcome the context window limitation, we propose using a context encoder to compress the original long contexts into a much more compact representation. The context encoder itself is a language model that takes a sequence of tokens as input and outputs a sequence of token embeddings. These output token embeddings, which we call summary embeddings, should be significantly shorter than the original context. The summary embeddings are then prepended to the LLM decoder and serve as the LLM’s system prompt. The user’s prompt is processed normally through the LLM decoder’s tokenizer, and the LLM generates answers (the assistant prompt) conditioned on the summary embeddings and user prompts.
In our design, the context encoder can be any model capable of producing a compact representation aligned with the LLM decoder. The summary embeddings can be viewed as pseudo-words in the LLM decoder’s text embedding space, representing abstract concepts or summaries. For our experiments, we select AutoCompressor for LLaMA2-7B (Chevalier et al., 2023), a context compressor finetuned on LLaMA2-7B with the dual ability to generate summary tokens from long contexts and perform text completions conditioned on the summary tokens. The compressor groups the document into chunks of 1536 tokens and recursively compresses each chunk to 50 summary embeddings. To ensure alignment between the summary embeddings and the decoder LLM, we choose LLaMA2-7B as our decoder LLM but use the AutoCompressor’s finetuned weights as the model weights.
While there are other available compressors for LLaMA2-7B (Ge et al., 2023; Mu et al., 2023), we find AutoCompressor to be most suited for our intended use cases given that it can 1) support compressing very long context due to its recursive training procedure, and 2) achieve a great compression ratio of 30:1. We consider the construction of context encoders (compressors) to be an important and orthogonal research direction. Developing more performant and universal context encoders in a streamlined fashion can further enhance the efficiency and effectiveness of LLoCO, denoting crucial future work.
3.2 Pipeline for Offline Context Learning
The pipeline of LLoCO consists of three primary stages: the preprocessing stage, the finetuning stage, and the serving stage. In this section, we introduce the pipeline and how to combine it with an existing RAG system to enable retrieval-augmented document QA with compressed context.
Preprocessing The preprocessing step in a traditional RAG system involves building a vector database to index a collection of documents. This process typically includes three main steps: 1) chunking the documents into passages if they exceed a certain length, 2) generating sentence embeddings of the passages using a text embedding model, and 3) indexing the passages using these embeddings. The result of the preprocessing step is a key-value map that associates each passage with its corresponding embedding.
In LLoCO, we introduce an additional preprocessing step that leverages our context encoder to process the original documents. Since we use AutoCompressor as our context encoder, we follow its practice of dividing the documents into chunks of 1536 tokens and passing them through the context encoder. The context encoder outputs 50 summary embeddings for each chunk recursively, with each embedding having a dimension of 4096. These summary token embeddings are stored alongside the original passages in the vector database, indexed by the sentence embedding of each passage.
Finetuning During the finetuning stage, we first segment the documents into groups based on their type (e.g., academic papers, news) or the tasks that users want to perform on them (e.g., QA, summarization). For each group of documents, we perform parameter-efficient finetuning using a LoRA adaptor. The finetuning data can be in-domain instruction pairs provided by the model provider. If such a dataset does not exist, it could also be generated using self-instruct (Wang et al., 2022) techniques or distilled from a more powerful model like GPT-4. At the end of the finetuning process, we obtain a finetuned LoRA adaptor for each group of documents. In the vector database, we store an identifier for the corresponding LoRA module alongside each passage entry. Additionally, we create a separate LoRA database to store all the LoRA adaptors.
Serving In a traditional RAG system, when a user asks a question, the system employs a retriever to fetch the top k relevant documents/passages from the vector database and concatenates them to the prompt to serve as the relevant context.
Figure 2: The serving stage of LLoCO’s pipeline.
In Figure 2, we illustrate LLoCO’s serving pipeline. Instead of retrieving the actual passages, we use the retriever to retrieve the compressed token embeddings of the passages. These token-embeddings are concatenated and prepended to the decoder LLM. Additionally, the system searches for the corresponding LoRA adaptor in the database and applies it to the decoder LLM. Applying a LoRA adaptor incurs minimal overhead to the system cost  (Hu et al., 2022). By leveraging recent work on serving multiple LoRA adaptors (Sheng et al., 2023a; Chen et al., 2023a), we can parallelize LoRA adaptor serving and simultaneously serve requests for documents associated with thousands of LoRA adaptors on a single GPU.
In our current system, we assume that all the passages retrieved from the vector database for a given user query belong to the same document group, allowing us to use a single dedicated LoRA adaptor. However, our approach can be extended to handle cases where the retrieved passages span multiple document groups. For instance, it is feasible to design algorithms to dynamically select the most relevant LoRA adaptor based on the majority of the retrieved documents or weight the contributions of different LoRA adaptors according to their relevance to the query (Muqeeth et al., 2024). Another interesting and orthogonal research direction is to explore composing multiple LoRA modules together (Huang et al., 2023), and we leave it as future work.
4 Experiments
In the experiment section, we aim to investigate the following aspects of LLoCO: (1) its effectiveness in comprehending compressed long contexts, (2) the extent to which summary embeddings can preserve essential information, and (3) the associated system costs.
Datasets We select four datasets dedicated to question-answering tasks—QuALITY (Pang et al., 2021), Qasper (Dasigi et al., 2021), NarrativeQA (Kočiskỳ et al., 2018), and HotpotQA (Yang et al., 2018)—along with one for summarization, QMSum (Zhong et al., 2021). All datasets have long documents as contexts. For all the datasets, we use their validation set for evaluation. We follow the official metrics for each dataset. For QuALITY, we report the exact match (EM) score. For QMSum, we report the geometric mean of ROUGE scores. For the remaining datasets (Qasper, NarrativeQA, and HotpotQA), we report the F1 scores. More details on these datasets can be found in Appendix A.2
Model Configuration In this study, we consider two base models. The first is the original LLaMA2-7B-chat (Touvron et al., 2023), which has a context window of 4096 tokens. The second is Longchat-7b-v1.5-32k (Li et al., 2023), a finetuned LLaMA2 model with an extended context window of 32,000 tokens via position interpolation. From now on, we will use LLaMA2-7B-4k to refer to the prior model and LLaMA2-7B-32k to denote the latter one. Unless otherwise specified, we set the LoRA rank to 8 for our experiments. All LLoCO’s models are finetuned on AutoCompressor, which is itself finetuned on LLaMA2-7B.
4.1 Long Document QA
To evaluate the effectiveness of LLoCO on the aforementioned long context datasets, we consider the following scenarios:
  1. 1.
    LLaMA-2-7B-4k/32k without Context. In this setting, we do not provide any documents to the LLM in the system prompt.
  2. 2.
    LLaMA-2-7B-4k/32k with Context. In this setting, we provide the LLMs with the ground truth document. We truncate the documents if their length exceeds the context window limit. This setting serves as the baseline.
  3. 3.
    LLaMA-2-7B-4k/32k with Retrieval. Retrieval is a standard baseline compression method for long document question answering. For each document, we chunk it into passages of 512 tokens and use Contriever (Izacard et al., 2021) to retrieve the top 5 passages from this document and concatenate them to form the context.
  4. 4.
    AutoCompressor. In this setting, we use AutoCompressor (Chevalier et al., 2023) to compress the contexts into summary embeddings and prepend them to the LLM, without performing any in-domain finetuning. This is equivalent to using an AutoCompressor straight out of the box. The AutoCompressor compresses a chunk of 1536 tokens into 50 summary tokens, resulting in an effective context window of roughly 128k tokens. We do not truncate the context unless it exceeds this limit.
  5. 5.
    LLoCO (ours). LLoCO is our proposed system for long document question answering. For each dataset we evaluate, we perform instruction finetuning using the question-answering pairs from the training set. During both finetuning and inference, we prepend the summary embeddings of the corresponding ground truth document to the LLM. We do not truncate the context unless it exceeds the 128k context window limit.
Our results are summarized in Table 1. When AutoCompressor is used out of the box without any in-domain finetuning, its performance sometimes falls short of the baseline where no context is appended. In contrast, by combining compression and finetuning, LLoCO significantly outperforms the baseline on all datasets by a substantial margin while using 30 times fewer tokens. Notably, for NarrativeQA, LLoCO achieves an F1 score of 28.34 compared to 14.42 for the baseline, showcasing the efficacy of our proposed technique.
Furthermore, the effectiveness of LLoCO is particularly evident in the case of the NarrativeQA dataset, where the average document length is 84,770 tokens, significantly exceeding the context window limit of LLaMA2-7B. By compressing the contexts into approximately 2,600 tokens on average, LLoCO is able to fit the compressed representations within the context window limit. This enables LLoCO to leverage all the available context, resulting in impressive F1 scores on this challenging dataset.
Setup Ctx Size $\tau$ QuA QAS QM NQA HQA Avg.
LLaMA2-7B
4k w.o. Context 4k 1x 32.16 7.68 12.73 10.85 22.22 17.13
32k w.o. Context 32k 1x 31.74 6.30 12.79 10.61 20.03 16.29
4k w. Context 4k 1x 40.45 16.67 14.62 14.42 32.47 23.44
32k w. Context 32k 1x 38.88 21.72 14.58 16.76 31.58 24.70
Compression Methods
LLaMA2-7B-4k w. Retrieval 4k 1.6x 38.73 18.29 14.33 22.28 27.95 24.31
LLaMA2-7B-32k w. Retrieval 32k 12.8x 36.48 24.92 15.40 19.32 22.32 23.68
AutoCompressor
128k 30x 33.51 15.03 12.53 11.66 21.01 18.75
LLoCO w. Full Context (ours) 128k 30x 41.51 29.01 16.68 28.34 37.83 30.67
Table 1: Experiment results on long document QA datasets. $\tau$ indicates the compression ratio. For LLaMA2-7B-4k/32k with retrieval, the compression ratio is obtained by dividing the model’s context window limit (4k/32k) by the length of the retrieved passages, which is fixed at 2560 tokens.
4.2 Ablation Study
Finetuning LLaMA with Original Context One interesting question is how well LoRA finetuning works over original uncompressed context, and how does that compare to LLoCO’s approach. To investigate this, we conduct additional finetuning experiments on LLaMA-7B-4k/32k, where we append the original uncompressed context as a system prompt during finetuning. These models are finetuned following the same setup as LLoCO.
As shown in Table 2, both LLaMA-7B-4k and LLaMA-7B-32k exhibit notable performance improvements after finetuning, with increases of $2.88\%$ and $6.62\%$ respectively. Despite these improvements, the performance of LLoCO remains comparable to that of the finetuned LLaMA-7B-32k model. This finding suggests that finetuning in a compressed context is just as effective as finetuning in the original context. Compared to full-context finetuning, LLoCO’s finetuning step is considerably faster and more cost-efficient due to the use of a significantly shorter context (see 4.5). This makes LLoCO a more practical finetuning solution for long document question-answering tasks.
Setup QuA QAS QM NQA HQA Avg.
LLaMA-7B-4k 40.45 16.67 14.62 14.42 32.47 23.44
LLaMA-7B-32k 38.88 21.72 14.58 16.76 31.58 24.70
LLaMA-7B-4k w. Finetuning 35.00 17.80 15.49 21.41 41.89 26.32 (+2.88%)
LLaMA-7B-32k w. Finetuning 40.12 29.71 16.36 28.72 41.68 31.32 (+6.62%)
LLoCO 41.51 29.01 16.68 28.34 37.83 30.67
Table 2: Comparison of LLoCO with LLaMA-7B baselines after in-domain finetuning.
Combined Instruction Finetuning In our previous experiments, we finetuned LLoCO on each dataset separately. In this ablation study, we investigate LLoCO’s performance when we combine all the training data and finetune a general model. To balance the dataset, we sample 10,000 examples from NarrativeQA and HotpotQA and use all examples from the other three datasets to form our final training set.
As shown in Table 3, LLoCO with combined finetuning surpasses baseline methods on all datasets except QMSum. The lower performance on QMSum can be attributed to the fact that it is a summarization task that favors longer, more detailed answers, but the combined finetuning process likely shifts the LLM’s behavior towards generating shorter, more focused answers, which may not be optimal for the summarization task.
Compared to in-domain finetuning, combined finetuning generally yields lower performance, with the exception of QuALITY. QuALITY questions heavily rely on the LLM’s reasoning abilities, whereas the other datasets primarily focus on the LLM’s capacity to retrieve relevant information from long contexts. We hypothesize that the additional finetuning on diverse datasets enhances the LLM’s ability to reason about long contexts, leading to improved performance on QuALITY. Overall, the combined finetuning approach demonstrates the potential for knowledge transfer across different tasks.
Setup QuA QAS QM NQA HQA Avg.
LLoCO w. Separate Finetuning 41.51 29.01 16.68 28.34 37.82 30.67
LLoCO w. Combined Finetuning 47.07 24.95 12.77 27.93 35.55 29.65
Table 3: Comparison of LLoCO with in-domain finetuning and combined finetuning.
4.3 Evaluation on LongBench
We further evaluate LLoCO on LongBench (Bai et al., 2023), which consists of 6 subtasks. Given that the primary applications of LLoCO are document question answering and summarization, our evaluation focuses on the SingleDoc QA, MultiDoc QA, and Summarization tasks. These three tasks overlap with some datasets (e.g. Qasper, QMSum) we evaluated in Section 4.1. However, the validation sets differ as LongBench samples a specific subset of examples for each dataset. We have rigorously ensured that LongBench does not feature any questions that overlap with those used in the LLoCO training, thereby completely removing any risk of data leakage. To achieve the best performance of LLoCO, we choose LoRA adaptors tailored for each particular category/dataset in our evaluation.
Specifically, for the SingleDoc tasks, we use the NQA LLoCO and Qasper LLoCo for the NQA and QAS tasks, respectively. For MultiField-En (MFE)  (Bai et al., 2023), we use GPT-4-turbo to generate question-answering pairs from the training documents and finetune LLoCO using these pairs. For MultiDoc tasks, we use the combined finetuned LLoCO from Section 2, which works well for all tasks. For the Summarization tasks, we create a training dataset by combining the entire QMSum (Zhong et al., 2021) training data with 5,000 randomly sampled examples from the MultiNews (Fabbri et al., 2019) training data. We then finetune LLoCO on this combined dataset.
SingleDoc MultiDoc Summarization
NQA QAS MFE HQA WMQA MSQ Gov QMS MNews Avg.
LLaMA-2-7B-4k 18.7 19.2 36.8 25.4 32.8 9.4 27.3 20.8 25.8 24.0
LLaMA-2-7B-32k 16.9 27.7 41.4 32.5 20.6 9.7 30.8 22.7 26.4 25.4
LLoCO 23.1 26.1 26.3 46.2 35.6 27.3 17.6 23.4 25.0 27.8
Table 4: Evaluation results on LongBench for SingleDoc, MultiDoc and Summarization. Numbers for LLaMA-2-7B-4k/32k are taken from the official LongBench’s repository (THUDM, 2023)
Our evaluation results show that LLoCO outperforms the LLaMA2 baseline on 5 out of the 9 datasets. In particular, LLoCO excels in the MultiDoc QA task, achieving significantly better results on all three datasets. LLoCO also demonstrates comparable performance in two datasets (MultiNews, Qasper), but falls short in the GovReport Huang et al. (2021) and MultiField-En datasets. The GovReport dataset is a summarization task that requires the model to generate a one-page summary. We found that LLoCO does not perform well when the generated content is lengthy, which is a limitation of our approach. The lower performance on the MultiField-En dataset could be attributed to the data being out-of-distribution compared to our training data, as this dataset does not provide any training examples. Despite these limitations, the average score of LLoCO across all datasets is higher than that of the LLaMA2 baseline, demonstrating the overall effectiveness of our approach.
4.4 Needle In A Haystack
We further investigate LLoCO’s proficiency to retrieve information across different positions of context window with varying lengths using the renowned Needle In A Haystack task (gkamradt, 2023). Tailoring this to our pipeline, we randomly select a long article exceeding 32k tokens from the NarrativeQA dataset as the “haystack”. This article is used to test the LLoCO model, which has been finetuned on this dataset as discussed in § 4.1.
[…]
(a) LLaMA2-7B-32k
(b) LLoCO
Figure 3: Fixed needle retrieval task. The sampled article (”haystack”) starts with ”Mary, ..., a gentle, fashionable girl...”, and a context-relevant needle was curated as ”Mary’s favorite fashion designer was Coco Chanel when she was a teenager. Who was Mary’s favorite fashion designer when she was a teenager?
[…]
(a) LLoCO (no finetuning)
(b) LLoCO (with finetuning)
Figure 4: Random needle retrieval with city-word pairs.
4.5 Inference Latency & Finetuning Throughput
In this section, We evaluate the inference latency improvements of our LLoCO method. The experiments are run on a single A100-80G-SXM GPU and a RTX A6000 GPU, both with a batch size of $1$ and a generation token count set to $16$. We measured the per-token latency with various context sizes. As illustrated in Figure 5, LLoCO realizes speed-ups of up to 7.62×7.62$\times$7.62 × on A100 and 7.19×7.19$\times$7.19 × on A6000 when compared to the LLaMA2-7B baseline without compression, under identical context conditions. While the baseline exceeds GPU VRAM limits for sequences longer than $32$k tokens, LLoCO maintains efficient generation for sequences up to $128$k tokens. Notably, LLoCO achieves the baseline’s $4$k latency for sequences that are 16$\times$ longer (64k) on the A100 and 32$\times$ longer (128k) on the A6000. Furthermore, LLoCO can process $32$k sequences on A6000 as fast as baseline’s $4$k on A100.
Additionally, we assess the finetuning throughput on the NarrativeQA dataset, which mainly consists of samples exceeding $32$k tokens. We compare the 7B $32$k baseline finetuned on 8 A100s and our LLoCO finetuned on both 8 A100s and 8 A6000s, all with a per-device batch size of 1 and a global batch size of 32, using 4 gradient accumulation steps. Figure 5 shows that our LLoCO achieves 11.52×11.52$\times$11.52 × and 6.82×6.82$\times$6.82 × throughput on A100s and A6000s, respectively. This highlights that LLoCO not only achieves competitive results, but also improves performance with much greater efficiency compared to finetuning the baseline with full context (as shown in § 4.2).
Figure 5: Left: End-to-end decoding per-token latency (ms) on A100 and A6000 GPUs. LLaMA2 without compression runs out of VRAM for sequences of 64k and 128k on A100, as well as for 32k sequences on A6000. Right: Train samples per second when finetuning on NarrativeQA for LLaMA2-7B-32k without compression and LLoCO.
5 Conclusion
We propose LLoCO, a new paradigm that addresses the long-context challenge by preprocessing the context before inference through parameter-efficient finetuning. Our approach extends the effective context window of a LLaMA2-7B model with a 4k tokens context size to handle up to 128k tokens. Evaluations on several long-context question-answering datasets show that LLoCO significantly outperforms in-context learning while using 30 times fewer tokens during inference, making it a promising solution for efficient long context processing.
Acknowledgement We thank Michael Luo, Sheng Shen, Chenglei Si and Siyuan Zhuang for their helpful comments and discussion. We also thank BAIR, Berkeley Deep Drive, Intel Corporation, and NVIDIA for supporting this research, as well as Hyperbolic Labs [^3] https://www.hyperbolic.xyz/ for providing the AI infrastructure for our experiments.
Appendix A Appendix
A.1 Extended Experimental Settings
In this part, we provide details of our experimental settings in the main text. For the experiments presented in Sec 4.1, we summarize the hyperparameter settings in Table 5. We use all samples from the train split of each dataset for finetuning, except for HotpotQA, from which we randomly select $20,000$ samples from its train split.
Hyperparameters QuA QAS QM NQA HQA
LoRA Rank $r$ 8
Optimizer AdamW
Weight Decay 0.00
[…]
LR Scheduler cosine annealing
Batch Size 8 8 8 32 32
Warmup Ratio 0.04
Epoch 3 3 3 1 3
Table 5: Hyperparameter settings of experiments.
A.2 More Details on Datasets
Here we provide more detailed descriptions of the 5 main datasets used for evaluation:
  • QuALITY (Pang et al., 2021) is a multiple-choice question-answering dataset over long contexts. It contains 150 articles with an average length of 5000 tokens, with 6737 questions in total. This dataset is particularly suited to evaluate models in a long-context setting.
  • Qasper (Dasigi et al., 2021) is a dataset for answering questions about NLP research papers, sourced from the Semantic Scholar Open Research Corpus. It includes various types of questions, ranging from detailed explanations to simple yes/no answers, based on provided documents.
  • QMSum (Zhong et al., 2021) features summaries and transcripts from meetings across various sectors, including academia and industry, focusing on query-based summarization. Participants are tasked with condensing dialogue transcripts based on specific queries.
  • NarrativeQA (Kočiskỳ et al., 2018) is a question-answering dataset derived from complete texts of books from Project Gutenberg and movie scripts from various sources. The challenge here involves generating a concise answer from long and potentially disorderly texts sourced from books.
  • HotpotQA (Yang et al., 2018)) is a Wikipedia-based dataset that challenges users to derive answers through multi-hop reasoning across several documents, featuring a broad range of questions beyond specific knowledge bases.